solve n(n+2) = 360 I am not sure how to do this one

Answer by rothauserc(4718)

rewrite the given equation by multiplying by n and regrouping we have
n^2 + 2n - 360 = 0
now we can use the quadratic equation
x = minus b plus or minus the square root of b^2 - 4ac divided by 2a
the general formula for solution is ax^2 + bx + c = 0
now we can substitute in the quadratic equation
x = -2 + square root(4-(4*1*-360)/2= 18
x = -2 - square root(4-(4*1*-360)/2= -20